### v1

# Chat with AI: The Surprising Turn of Real-time Video Communication from Human to AI

Abstract.
AI Video Chat emerges as a new paradigm for Real-time Communication (RTC), where one peer is not a human, but a Multimodal Large Language Model (MLLM). This makes interaction between humans and AI more intuitive, as if chatting face-to-face with a real person. However, this poses significant challenges to latency, because the MLLM inference takes up most of the response time, leaving very little time for video streaming. Due to network uncertainty and instability, transmission latency becomes a critical bottleneck preventing AI from being like a real person. To address this, we propose Artic, an AI-oriented Real-time Communication framework, exploring the network requirement shift from ”humans watching video” to ”AI understanding video”. To reduce bitrate dramatically while maintaining MLLM accuracy, we propose Context-Aware Video Streaming that recognizes the importance of each video region for chat and allocates bitrate almost exclusively to chat-important regions. To avoid packet retransmission, we propose Loss-Resilient Adaptive Frame Rate that leverages previous frames to substitute for lost/delayed frames while avoiding bitrate waste. To evaluate the impact of video streaming quality on MLLM accuracy, we build the first benchmark, named Degraded Video Understanding Benchmark (DeViBench). Finally, we discuss some open questions and ongoing solutions for AI Video Chat.
1. Introduction
AI Video Chat is a new paradigm for Real-time Communication (RTC). Since OpenAI released GPT-4o (Hurst et al., 2024), Multimodal Large Language Models (MLLMs) have been continuously emerging, such as Qwen2.5-Omni (Xu et al., 2025), Gemini-1.5 (Team et al., 2024), VITA-1.5 (Fu et al., 2025), and OmniLive (Zhang et al., 2024). Compared to traditional LLMs, MLLMs enable users to directly input video and audio for interaction, rather than just text. This makes the interaction between humans and AI more intuitive as if chatting face-to-face with a real person (Chen et al., 2024b). However, MLLMs require high-performance computing devices to support real-time inference (such as 8*A100 GPUs (Qiu et al., [n. d.])). Mobile devices (like phones or smart glasses) cannot meet the computing requirements, which makes MLLMs inevitably deployed in the cloud. So in existing systems, the client sends user video and audio to the cloud for MLLM inference, and the cloud then feeds back responses to users, as shown in Figure 1.
Figure 1. AI Video Chat is a new paradigm for real-time communication. The user sends video and audio to the AI for thinking. The AI feeds back to the user. Low latency is crucial for making AI act like a real person.
AI Video Chat raises significant challenges to RTC transmission latency. To ensure a fluent interactive experience, the end-to-end response latency of video chat needs to remain below 300 ms (Lai et al., 2022). In traditional video chat, the human peer on the other side can respond instantly. Therefore, transmission latency accounts for the vast majority of end-to-end response latency. To reduce transmission latency, current state-of-the-art RTC frameworks (such as WebRTC (Carlucci et al., 2016)) adopt technologies like Adaptive Bitrate (ABR) (Mao et al., 2017; Yan et al., 2020; Akhtar et al., 2018; Huang et al., 2014; Jiang et al., 2012), Congestion Control (Cardwell et al., 2017; Dong et al., 2015, 2018; Meng and Xu, 2024), and Forward Error Correction (FEC) (Meng et al., 2024; An et al., 2025; Rudow et al., 2023; Cheng et al., 2024; Li et al., 2023) to satisfy user experience. However, in AI Video Chat, responses are generated through MLLM in an autoregressive manner, which is time-consuming. Even when inputting only audio tokens, the computational latency is at least 232 ms (Hurst et al., 2024). To constrain the response latency below 300 ms, the time left for transmission is at most 68 ms, which is difficult to guarantee. So the large latency makes users clearly feel that the other side is not a real person.
Is it possible to reduce the latency of AI Video Chat to an extremely low level? This vision allows us to grasp the ”holy grail” (Chen et al., 2024b) of AI research from the perspective of network systems: making AI like real humans.
To realize this vision, we propose Artic, an AI-oriented Real-time Communication framework, exploring the network requirement shift from ”humans watching video” to ”AI understanding video”. We begin by recognizing the main differences between AI Video Chat and traditional RTC: First, QoE changes from human perceptual quality to MLLM response accuracy. Second, the receiver throughput is far lower than the sender throughput. Third, uplink is more pressing than Downlink. Then, through prototype measurements, we identify two key factors for low latency: ultra-low bitrate and loss resilience. Based on three insights, we make contributions:
  • Video should be Context-Aware (§3.2). To reduce bitrate dramatically while maintaining MLLM accuracy, we propose Context-Aware Video Streaming, allocating more bitrate to chat-important video regions while allocating as little bitrate as possible to chat-irrelevant regions.
  • Frame Rate is FEC (§3.3). Since MLLMs process videos at a very low frame rate, most received frames are redundant. Although redundant frames cause bitrate waste, they can also substitute for lost/delayed frames. So we propose Loss-Resilient Adaptive Frame Rate to simultaneously minimize bitrate waste and packet retransmission.
  • The First Benchmark (§3.1). Considering that there is no benchmark that can evaluate the impact of video quality on MLLM accuracy, we propose the first one, named Degraded Video Understanding Benchmark (DeViBench).
2. Motivation
2.1. Main differences between AI Video Chat and traditional RTC
QoE changes from human perceptual quality to MLLM response accuracy. In traditional RTC, QoE focuses on measuring the perceptual quality of human eyes. For example, using penalty terms like stalling time (Cheng et al., 2024; Yan et al., 2020) or quality variance (Chen et al., 2024a) to measure temporal stability. Using SSIM (Yan et al., 2020) or VMAF (vma, 2025) to measure visual quality. However, in AI Video Chat, the viewer of the video changes from humans to MLLMs. At this point, most perception-related metrics are no longer needed, and the optimization objective of QoE becomes the accuracy and latency of MLLM’s responses. So RTC strategy can undergo a significant turn. For example, to reduce latency, temporal stability (such as frequent bitrate adjustments) and visual quality (like lowering bitrate in Figure 3) can be sacrificed, as long as accuracy is enough.
The receiver throughput is far lower than the sender throughput. In traditional RTC, the data throughput at the receiver is comparable to that at the sender, for example, both are 1920*1080 resolution at a 30 FPS frame rate. However, in AI Video Chat, MLLM is limited by context length (finite number of tokens) and cannot fully process the received video. Therefore, the received video needs to be actively downsampled before being fed to the MLLM. In terms of frame rate, existing AI Video Chat systems support a maximum processing rate of only 2 FPS (Xu et al., 2025; Wang et al., 2025b; Fu et al., 2025). In terms of resolution, regardless of how high the original resolution is, it will be downsampled to no more than 602,112 pixels (Xu et al., 2025).
Uplink is more pressing than Downlink. In traditional RTC, each peer is both video sender and receiver. In contrast, AI Video Chat is unidirectional video transmission, where the user only acts as the video sender and MLLM only acts as the video receiver. MLLM sends responses to the user in the form of audio or text, and these representations have much lower bitrates than video. Thus, uplink needs better network conditions than downlink, for example, larger bandwidth.
2.2. What factors affect the transmission latency of AI Video Chat?
Figure 2. How bitrate and packet loss affect latency (with 10 Mbps bandwidth). To optimize video quality, traditional RTC systems select bitrate from the gray region. But in AI video chat, to maintain accuracy, we only need to select bitrate from the yellow region (§2.2).
To analyze the factors affecting transmission latency in AI Video Chat, we build a prototype and conduct preliminary measurements. Specifically, we implement a WebRTC-based unidirectional video transmission system and a network emulator. Under given bandwidth (10 Mbps) and one-way network delay (30 ms), we run video transmission for a total duration of 40,489 seconds, and collect statistics on frame latency (the time difference from the frame being sent to being completely received) with different packet loss rates and bitrates, as shown in Figure 2. The results indicate:
First, when the bitrate exceeds the bandwidth, frame latency becomes enormous. This is because excessive bitrate causes congestion, where packet accumulation causes latency to increase rapidly. Therefore, existing RTC systems employ ABR algorithms to set the bitrate as close as possible to (but below) the bandwidth, maximizing video quality while avoiding stalling, as shown in the grey region of Figure 2. Second, when the bitrate does not exceed the bandwidth, frame latency also increases as the bitrate increases. This is due to the fact that each packet has a limited size (around 1400 bytes). A higher bitrate means each frame will be divided into more packets. Due to packet loss, more packets mean the probability of each frame being completely received in one attempt decreases. For packets that are not received, retransmission may be required, leading to increased latency. Therefore, even when the bitrate is below the bandwidth, AI Video Chat can further reduce the bitrate to achieve lower latency. This differs from traditional ABR and offers another space for bitrate selection, as shown in the yellow region of Figure 2. Third, when the packet loss rate increases, frame latency increases. Similarly, this is also caused by more retransmissions.
2.3. Key Insights and Potential Gains
Video should be Context-Aware. According to §2.2 reducing video bitrate can decrease transmission latency. To reduce bitrate, existing methods typically increase quantization parameters or lower resolution (Schwarz et al., 2007), which inevitably degrades video quality. Interestingly, the degradation in video quality does not necessarily lead to a decrease in MLLM accuracy, which depends on the current chat context. As illustrated in Figure 3, when the user asks ”Could you tell me the present score of the game?”, even if the video bitrate is reduced from 4000 Kbps to 200 Kbps, the MLLM can still answer accurately. However, when the user asks ”What logo is seen on the jersey of the player covering his mouth?”, the blurry video leads to incorrect responses. This is because, in different chat contexts, the MLLM needs to focus on different video regions. Meanwhile, different video regions are affected differently by low bitrate. Thus, rather than reducing bitrate in a context-agnostic manner, the video should be context-aware. More bitrate should be allocated to chat-important regions, while less bitrate should be allocated to chat-irrelevant regions.
Figure 3. Why video should be context-aware in AI Video Chat. In the first dialogue, even if the video bitrate decreases from 4000 Kbps to 200 Kbps, the MLLM can still response accurately. But in the second dialogue, the blurry video leads to incorrect responses. Therefore, rather than reducing bitrate in a context-agnostic manner, bitrate allocation should be determined by the current chat context (§2.3).
How to be aware of the chat context? Our idea is: the user words can indicate which video regions are important for the current chat. Therefore, we can take the user words as a reference to compute the semantic correlation of different video regions. For this, we adopt the Contrastive Language-Image Pre-Training (CLIP) model (Radford et al., 2021), which maps images and language to the same feature space. Hence, to derive semantic correlation, we only need to compute the similarity of features between video regions and user words. We show some examples in Figure 5, which demonstrates that user words and CLIP can well point out the importance of different video regions for chatting. For example, when the user asks ”Is the dog in the video erect-eared or floppy-eared?”, the dog’s head region exhibits the highest correlation. On the other hand, even when the user words do not explicitly indicate the object, CLIP can still estimate correlation based on high-level understanding. For example, when the user asks ”Infer what season it might be in the video”, grass has the highest correlation. This is because the growth of grass can imply the current season (CLIP even ignores the blurry grass in the distance). Thus, this context-aware mechanism allows us to optimally allocate the bitrate (§3.2).
Figure 4. Frame Rate is FEC. MLLM processes video at a very low frame rate (green region), so higher frame rates will introduce redundancy (red region). Higher frame rates on one hand cause bitrate waste, while on the other hand enhance loss resilience (§2.3).
Figure 5. How to achieve context awareness? The user words can indicate which regions in the video are important for the current chat context. Based on CLIP, we can even recognize important regions through high-level understanding. For example, in the third dialogue, the growth of grass implies the current season (§2.3).
Frame Rate is FEC. According to §2.2, besides bitrate, packet loss rate also has an important impact on latency. We find that frame rate can make trade-offs between bitrate and packet loss rate in existing AI Video Chat systems. As described in §2.1, MLLM processes videos at very low frame rates (such as 2 FPS), while the client still transmits videos at conventional frame rates (like 30 FPS). After receiving the video, MLLM will sample frames at a uniform time interval (such as 500 ms) for processing, while the remaining frames are not used. This causes most received video frames to be redundant. So to reduce video bitrate, the most reasonable solution is to use the same low frame rate (2 FPS) at the sender. Figure 4 shows the relationship between frame rate and video bitrate. It can be seen that lowering the frame rate can reduce bitrate while not affecting MLLM accuracy.
Figure 6. How to achieve loss resilience? When the expected frame is not completely received on time, the MLLM directly discards it and uses the previous redundant frame, without waiting for retransmission.
In contrast to reducing bitrate, increasing the frame rate can enhance resilience to packet loss. As illustrated in Figure 6, rather than sending video at 2 FPS, the client can send at a higher frame rate. When the expected frame (e.g., sampled at 500 ms intervals) is not received on time, the MLLM can directly discard it and use a previous (redundant) frame, without waiting for retransmission. In this way, redundant frames can handle packet loss in an FEC manner. This presents a trade-off: higher frame rates increase redundancy for better loss tolerance, but also lead to higher bitrates. So the frame rate should be selected based on loss rate (§3.3).
The first benchmark evaluates how video streaming quality affects MLLM accuracy. According to §2.1, QoE metrics in AI Video Chat change from perception to accuracy. This causes existing benchmarks in the video streaming field to be inapplicable, as they focus on perceptual quality and do not involve response accuracy. In the MLLM field, there are some benchmarks targeting Streaming Video Understanding tasks (Wang et al., 2025b; Yang et al., [n. d.]), such as StreamingBench (Lin et al., 2024). In these benchmarks, each video includes several Question-Answer (QA) samples for evaluating the response accuracy. However, these benchmarks aim to test the MLLM’s intelligence, so all the input videos are ideally high-bitrate (e.g., 4000 Kbps).
To evaluate how video streaming quality affects accuracy, we transcode videos from StreamingBench to 200 Kbps. Then we conduct testing on these low-bitrate videos with the original QA samples. The results show that only 8% of QA samples are answered incorrectly at low bitrate and correctly at high bitrate. This is because the QA samples in StreamingBench are too simple and high-level, requiring only coarse-grained video content to answer correctly. For example, in Figure 3, when the question is ”What is the player doing?”, even if the video quality is particularly poor, the MLLM can still provide the correct answer ”He is covering his mouth.” However, in real-world scenarios, there are often many detail-rich questions that are very sensitive to video quality. For example, in Figure 3, when the question is ”How many spectators can be seen?”, even slight blurriness will prevent the MLLM from providing the correct answer. So it is necessary to establish a more challenging benchmark to reflect the real-world impact of video degradation on MLLM accuracy (§3.1).
3. Towards RTC for AI
For the above insights, we begin by constructing the first benchmark evaluating how video quality affects MLLM accuracy, named Degraded Video Understanding Benchmark (DeViBench). Then we design two modules: context-aware video streaming and loss-resilient adaptive frame Rate.
3.1. DeViBench
Figure 7. DeViBench’s pipeline for automatic QA sample construction. Details can be found in §3.1.
In this section, we propose DeViBench. As described in §2.3, we need to construct QA samples that are sensitive to video quality. For this, the most straightforward way is to hire volunteers to ask tricky questions about degraded videos. However, this is too expensive and inefficient, hindering the scale-up of the dataset. So we ask: Can QA samples be constructed automatically and cheaply? Rethinking the background of AI Video Chat, MLLMs are already capable of understanding videos and giving responses. So we leverage MLLMs to replace human volunteers and develop an automatic QA sample construction pipeline. As illustrated in Figure 7, this pipeline consists of 5 steps:
Video Collection. We first collect videos to ask questions. To align with the domain and scale of existing MLLM benchmarks (Lin et al., 2024), we directly use their videos (discarding QA).
Video Preprocessing. To allow MLLMs to understand the quality degradation caused by low bitrates, we transcode the original videos to low-bitrate versions (200 Kbps). The low bitrate video and the original video are horizontally concatenated into one video. Then this concatenated video will be input to the MLLM for understanding and QA generation.
QA Generation. To enable MLLMs to generate QA pairs based on the concatenated video, we carefully designed a prompt with guidance from persona, context, core task, execution steps, constraints, and output format, as shown in Figure 8. This prompt ensures that MLLMs can recognize quality differences and generate quality-sensitive QA pairs.
QA Filtering. The generated QA pairs will be filtered. We separately input the original video and the low bitrate video into the MLLM and use the generated QA pairs for questioning. If the answer from the original video is correct and the answer from the low bitrate video is wrong, we accept this QA pair. In practice, 25.2% of the QA pairs can be accepted.
Cross Verification. Since the answer generated by MLLM may also be incorrect, this cannot be filtered out through the above testing. Hence, we utilize another MLLM for cross-verification. We feed the above accepted question into another MLLM, and if the new answer is consistent with the above accepted answer, we finally approve this QA pair. In our experiments, 57.7% of the accepted QA pairs can pass cross-verification. Considering all the above validations together, finally 14.6% of the generated QA pairs are valid.
Figure 8. Our prompt for QA Sample Generation.
3.2. Context-Aware Video Streaming
In this section, we describe how to achieve context-aware streaming, significantly reducing bitrate while maintaining MLLM accuracy. According to §2.3, we first leverage the CLIP model to compute the semantic correlation between user words and video regions. To ensure real-time computing on mobile devices, we adopt Mobile-Clip (Vasu et al., 2024). Specifically, given the current user words $\mathcal{T}$ and the latest video frame $F\in\mathbb{R}^{H\times W\times 3}$, we first partition $F$ into non-overlapping patches $\left\{P_{mn}\mid 1\leq m\leq\left\lfloor H/N\right\rfloor,1\leq n\leq\left\lfloor W/N\right\rfloor\right\}$, where each patch $P_{mn}\in\mathbb{R}^{N\times N\times 3}$ represents a video region. Then, the CLIP visual encoder $\phi_{v}(\cdot):\mathbb{R}^{N\times N\times 3}\rightarrow\mathbb{R}^{d}$ is employed to extract patch-wise features $f_{mn}^{v}=\phi_{v}\left(P_{mn}\right)$, while the CLIP language encoder $\phi_{l}(\cdot):\mathcal{T}\rightarrow\mathbb{R}^{d}$ encodes the user words $\mathcal{T}$ into a semantic features $f^{l}=\phi_{l}(\mathcal{T})$. Here $d$ denotes the unified feature dimension. Semantic correlation $\rho_{mn}$ between user words and patches are then computed as cosine similarities:
(1) $\rho_{mn}=\frac{f_{mn}^{v}\cdot f^{l}}{\left\|f_{mn}^{v}\right\|\left\|f^{l}\right\|}\in[-1,1]$
Semantic correlation $\rho_{mn}$ can measure the importance of region $P_{mn}$ for the current chat context. The larger $\rho_{mn}$ is, the more important $P_{mn}$ is. So we can allocate more bitrate to important regions while allocating as little bitrate as possible to irrelevant regions. To achieve this, we adjust the Quantization Parameters (QP) of different regions during video encoding. When QP is larger ($0\leq\mathrm{QP}\leq$ 51), the region occupies less bitrate, but the quality becomes worse. Specifically, for region $P_{mn}$, its $\mathrm{QP}_{mn}$ is derived as:
(2) $\mathrm{QP}_{mn}=51\left(1-\left(\frac{\rho_{mn}+1}{2}\right)^{\gamma}\right)$
Where $\gamma$ is the temperature parameter, which we set 3 to aggressively penalizes irrelevant regions ($\rho_{mn}\ll 1$).
Figure 9. Context-aware streaming can dramatically lower the bitrate while maintaining MLLM accuracy.
We evaluate the performance gains in Figure 9. The results demonstrate that context-aware streaming can dramatically lower the bitrate while maintaining MLLM accuracy. For example, when the bitrate is reduced from 800 Kbps to 400 Kbps (50% reduction), the MLLM (Qwen2.5-Omni (Xu et al., 2025)) accuracy drops from 0.73 to 0.33. After integrating context-aware streaming, the accuracy only decreases from 0.93 to 0.87.
3.3. Loss-Resilient Adaptive Frame Rate
In this section, we describe how to select the frame rate based on the current loss rate, thereby jointly reducing redundancy and packet retransmission. According to Figure 6, when the expected frame is not received, Artic uses a previous redundant frame as the substitute. So given the loss rate, we aim to minimize the frame rate while maximizing the probability that at least one redundant frame is received on time.
Specifically, let $p_{t}\in[0,1]$ denote the packet loss rate during decision epoch $t$ (simplified to i.i.d. when the decision epoch is sufficiently small). Let $\kappa_{t}\in\mathbb{Z}^{+}$represent the average number of packets per frame within epoch $t$. The probability of successfully receiving a frame is then given by:
(3) $P_{f}=\left(1-p_{t}\right)^{\kappa_{t}}.$
Let $R_{M}$ denote the inference frame rate of the MLLM and let $R_{t}$ represent the video frame rate. So the MLLM samples one frame from every $K=$ $\left\lceil R_{t}/R_{M}\right\rceil$ frames. The probability of successfully receiving at least one frame in $K$ frames is:
(4) $P_{s}=1-(1-P_{f})^{K}$
To avoid retransmissions, it should satisfy $P_{s}\geq(1-\epsilon)$, where $\epsilon=0.001$ is the residual term. Therefore, we finally select the minimum frame rate $R_{t}$:
(5) $R_{t}\geq R_{M}\cdot\left\lceil\frac{\ln\epsilon}{\ln\left(1-\left(1-p_{t}\right)^{\kappa_{t}}\right)}\right\rceil$
We evaluate the performance gains in Figure 10. The preliminary results show that our method can reduce the latency caused by stalling from 25.3 ms to 1.2 ms, while avoiding the bitrate waste brought by high frame rates.
Figure 10. Eliminating the latency caused by stalling, while avoiding the waste brought by high frame rates.
4. DISCUSSIONS AND OPEN QUESTIONS
Proactive context-aware. In this paper, Artic leverages user words to achieve context awareness. It may not necessarily perform well in practice. Because it requires user words to be known before video encoding. But users may speak at any moment in the video, causing user words not to cover some segments. For example, in some benchmarks like (Lin et al., 2024; Wang et al., 2025b; Yang et al., [n. d.]), they assume that users ask questions at the end of the video. As our next step, we are building a proactive context-aware mechanism that can actively recognize important video regions even if users do not speak.
MLLM long-term memory. To minimize bitrate, Artic discards most video content irrelevant to the current chat context. This is based on the assumption that the current chat only references real-time video content. However, some MLLMs have developed long-term memory mechanisms (Wang et al., 2025a; Xiong et al., 2025; Qian et al., 2025), allowing chats to reference historical video content. Some video content, even if not relevant in the current chat context, may be needed in future chats. As our next step, we are developing a semantic layered video streaming framework. Different from SVC (Schwarz et al., 2007) that layers based on video quality, we layer by semantic correlation. The base layer contains the most important video content for the current chat context, so it must ensure low latency. The enhancement layers contain complete video details, used to offline build long-term memory, so they are not sensitive to latency.
Token pruning. To further lower the end-to-end latency, it is necessary to reduce the inference latency of MLLMs. Since MLLMs run in an autoregressive manner, a straightforward solution is to decrease the number of input tokens. Some related work exploits attention mechanisms (Zhong et al., 2024) or video redundancy (Yao et al., 2025) to prune most visual tokens, without affecting MLLM accuracy. In this paper, context-aware streaming has already recognized important video regions, so it makes much sense to prune tokens from chat-irrelevant regions. As our next step, we are developing context-aware token pruning mechanisms to accelerate MLLM inference.